Find the Laplace Transform of half-wave and full-wave

rectified sine wave given in the following figures.

Take w = 2


1
Expert's answer
2021-05-24T18:57:44-0400

Laplace transform of half wave rectified sine wave is given by -


f(t)= { "sin(\omega t)" , 0 < t < "\dfrac{\pi}{\omega}"

{ 0 , "\dfrac{\pi}{\omega}" < t < "\dfrac{2{\pi}}{\omega}"


we know that the time period of half wave rectifier is given by "T=\dfrac{2{\pi}}{w}"



If we know the time period of given wave function , then laplace transfrom can be calculated by using formula which is given by


"L[{f(t)}]=\dfrac{1}{1-e^{(-sT)}}\int_{0}^{T}e^{-st}f(t)dt"




"L[{f(t)}]=\dfrac{1}{1-e^{(-s\dfrac {2{\pi}}{\omega})}}\int_{0}^{T}e^{-st}f(t)dt"


Now , we will split this integeral in the domain which is given above ,


"L[{f(t)}]=\dfrac{1}{1-e^{(-sT)}}[\int_{0}^{\dfrac {{\pi}}{\omega}}e^{-sT}sin({\omega}t)dt" +"\int_{\dfrac{\pi}{\omega}}^{\dfrac {2{\pi}}{\omega}}e^{-sT}0 \ dt" ].....1)


"L[{f(t)}]=\dfrac{1}{1-e^{(-sT)}}[\int_{0}^{\dfrac{\pi}{\omega}}e^{-sT}sin({\omega}t)dt]"


We know integeration of "\int{e^{at}sin(bt)}" ="\dfrac{e^{at}}{a^{2}+b^{2}}[asin(bt)-bcos(bt)]"


now applying the above result in , in the integeral "\int_{0}^{\dfrac{\pi}{\omega}}e^{-sT}sin(\omega t)dt"


"=\dfrac{e^{-st}}{s^{2}+({\omega })^{2}}[-ssin{\omega}t-{\omega}cos{\omega}t]_{0}^{\dfrac{\pi}{\omega}}"


"=" now putting the limits , in above integeral , we conclude ,


"=\dfrac{e^{-s\dfrac{\pi}{\omega}}}{s^{2}+({\omega })^{2}}[{-ssin{\pi}-{\omega}cos{\pi}}]-\dfrac{e^{0}}{s^{2}+({\omega })^{2}}[{0-{\omega}cos0}]"


="\dfrac {\omega}{s^{2}+{\omega}^{2}}" ("1+e^{\dfrac{-s{\pi}}{\omega}}" )........2



now putting these integeration result in equation....... 1 ) we get -



"\dfrac{1}{1-e^{\dfrac{-2s{\pi}}{\omega}}}" ["\dfrac{\omega}{s^{2}+{\omega}^{2}}(1+e^{\dfrac{s{\pi}}{\omega}})" ]




This is the laplace transfrom of half wave rectifier .



Laplace transform of full wave rectifier sin equation is -


f(t)= { "sin(\omega t)" , 0 < t < "\dfrac{\pi}{\omega}"



The period of function is "\dfrac{\pi}{\omega}"



"L[{f(t)}]=\dfrac{1}{1-e^{(-sT)}}\int_{0}^{T}e^{-st}f(t)dt"



"L[{f(t)}]=\dfrac{1}{1-e^{(-sT)}}\int_{0}^{T}e^{-st}sin{\omega}tdt"



"L[{f(t)}]=\dfrac{1}{1-e^{(-sT)}}[\int_{0}^{\dfrac {{\pi}}{\omega}}e^{-sT}sin({\omega}t)dt]"




= "\dfrac{1}{1-e^{\dfrac{-2s{\pi}}{\omega}}}[\dfrac{\omega}{s^{2}+{\omega}^{2}}(1+e^{\dfrac{s{\pi}}{\omega}})]"